Please help me solve this: Lawrence has a daughter, Janet. Five years ago, he was 4 times as old as Janet. Two years from now, he
will be three times as old as Janet. How old is Lawrence now?

Answer by math_helper(2461)


L = Lawrence's age
J = Janet's age

L - 4 = 4 (J - 4) four years ago
L + 2 = 3 (J + 2) two years from now
L = 4J-12
L = 3J+4
Subtracting bottom eq from top eq:
0 = J+8 ==> J=16 ==> L=52
Ans: Lawrence is 52
Check: Four years ago: Lawrence was 48, Janet was 12, 48=4*12 (ok)
Two years from now: Lawrence will be 54, Janet will be 18, 54=3*18 (ok)